Exercise 15-24 Sage Inc. began operations in January 2015 and reported the following resuits for each of its 3 years of operations. 2015 $277,000 net loss 2016 $40,000 net loss 2017 $812,000 net income At December 31, 2017, Sage Inc. capital accounts were as follows 8% cumulative preferred stock, par value s 100; authorized, issued, $550,000 and outstanding 5,500 shares Common stock, par value $1.00; authorized 1,000,000 shares issued and outstanding 708,000 shares $708,000 Sage Inc. has never paid a cash or stock dividend. There has been no change in the captal accounts since Sage began operations. The state law permits dividends only from retained earnings (e) Compute the book value of the common stock at December 31, 2017. (Round answers to 2 decimal places, e.g. $38.5o. Book value per share b) Compute the book value of the common stock at December 31, 2017, assuming that the prefered stock has a lgu dating velue of Sİ06 per share. (eo ed as to ent places, e.g $38.50. Book value per share Question Attempts: 0 of 3 used SAVE FOR LATER SUBHIT ANSWER Rights Reserved. A Division of Version 4.24.1
Answer #1

Solution a:

Retained earnings on December 31, 2017 = $812,000 - $277,000 - $40,000 = $495,000

Retained earnings for common shareholders = Total retained earnings - Preferred dividend for 3 years

= $495,000 - ($550,000*8%*3) = $363,000

Book value per share of common stock = ($708,000 + $363,000) / 708000 = $1.51 per share

Solution b:

Book value per share = ($708,000 + $363,000 - $550,000*6%) / 708000 = $1.47 per share